• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name brads

    Log entry time 22:29:26 on February 07, 2009

    Entry number 259586

    This entry is a followup to: 259554

    keyword=BB Cerenkov work: TDC delay cables fixed, Disc. output widths increased

    The 360ns ribbon cable delays that were added yesterday had a few problems. There were some cables with a polarity flip from end-to-end and a couple had shorts (at least when I looked at them). I expect this was the main reason for the really weird looking TDC plots...

    I also took a look at the output widths on all of the Cerenkov related discriminators and set them to 100ns. Most had been set to 40, but there was one that was only 20ns. There is some high frequency ringing that might have been causing some multiple hits with the old shorter pulses.

    PMT 12 is now showing the high-frequency (~5ns period) bipolar noise on the pulse that was seen on PMT 2 yesterday. It just occurred to me that maybe PMT 12 was confused with PMT 2 last night -- I foolishly didn't check the analog signal of PMT 2 this afternoon. I did check the TDC signal for PMT 2 and it wasn't too bad.